How many different two-digit numbers can be formed using the digits 1, 3, 4, and 5. Repetition of digits is allowed.

How many 7-digit phone numbers can be formed if the first digit cannot be 0 and repetition of digits is allowed?

Answer by stanbon(48558)

How many different two-digit numbers can be formed using the digits 1, 3, 4, and 5. Repetition of digits is allowed.
------
4*4 = 16 two-digit numbers
=================
How many 7-digit phone numbers can be formed if the first digit cannot be 0 and repetition of digits is allowed?
-------
ways to choose the first digit = 9
ways to choose remaining 6 digits = 10^6
-------
Total number of 7-digit phone numbers = 9*10^6 = 9 million
=============
Cheers,
Stan H.